A cylinder of diameter d and length L is required to float in a liquid. How much L for the cylinder to float when its longitudinal axis being vertical? (specific gravity for cylinder and liquid are 0.7 & 0.9 respectively)


1
Expert's answer
2021-04-01T01:09:31-0400

Let the length of cylinder (H) is submerged in the liquid when it is kept in the liquid such that it's longitudinal axis is vertical.

Therefore in stable equilibrium condition,

The weight of liquid displaced=weight of cylinder

If: d=diameter of cylinder

L=length of cylinder

Sc= specific gravity of the cylinder

SL= specific gravity of liquid

Therefore,

Weight of liquid displaced=Weight of cylinder

"S_L× \frac{π}{4}×d^2×H= S_c × \frac{π}{4}×d^2×L"

Hence,

"H= \frac{S_c}{S_L}×L=\frac{0.7}{0.9}×L=\frac{7}{9}×L"

Therefore length of cylinder submerged in the liquid is,

H="\frac{7}{9}×L"

Therefore the length of cylinder which is floating outside the liquid.

=L- H="\frac{2}{9}×L=0.222×L"